OpenStudy (anonymous):

Find the sum 1 2/5 + 2 1/10=?

OpenStudy (anonymous):

improper fractions are just wrong lol

OpenStudy (anonymous):

Lol helppp D:

hero (hero):

1.4 +2.1 = 3.5 = 3 1/2 Easy

OpenStudy (anonymous):

10/5 + 21/10

OpenStudy (anonymous):

Well arn't you the smart cookie in the jar!! :P

OpenStudy (anonymous):

20/10 + 21/10 = 41/10 = 4 1/10 but I could be wrong

OpenStudy (anonymous):

yep, I'm wrong lol